# Dynamic tariffs-based demand response in retail electricity market under uncertainty

### v5

@@ 4.5 Market outcomes: Electricity prices and retailer’s quantities comparison
Let’s expand the comparisons to include price tariffs, quantities, consumers’ utility parameters, and demand flexibility. Figure 5 presents the price tariffs sent by the retailer and quantities purchased by the retailer from the spot market in both models. Price tariffs are plotted on the left axis, and spot market quantities on the right axis. This allows for an analysis of how changes in utility parameters impact spot market purchases and price tariffs. Figure 5(a) depicts the Benchmark case, serving as a baseline for comparing the impacts of other parameters. In the quadratic utility function, the linear term (as seen in Figure 5 (b)) captures the initial positive impact of electricity consumption on consumers’ utility. Meanwhile, the Quadratic case models the diminishing returns in satisfaction as electricity consumption increases. As a result, when parameter $B_{jt\omega}$ increases, the retailer purchases a larger quantity from the wholesale market, as illustrated in Figure 5(c), and subsequently sells it back to consumers at lower price tariffs in the retail market. The lower price, in turn, increases consumer utility. It should be noted that from the retailer’s perspective, an increase in $B_{jt\omega}$ has an effect analogous to an increase in demand from the consumers.
Overall, the electricity price tariffs sent by the retailer and the quantities bought by the retailer from the spot market exhibit a similar trend. The proposed models effectively characterize the behaviors of market participants, facilitating the generalization of managerial and economic insights. Under the MILP model, consumers experience increased welfare due to lower prices compared to the MPEC model. Conversely, the MPEC model benefits the retailer, who acts as a price-maker, by enhancing its expected profits. Additionally, while prices remain relatively stable in the MPEC model, the quantities purchased from the spot market are significantly affected by the flexibility of consumers. Most notably, high consumer flexibility significantly impacts the quantity of purchases in the spot market. There is an uptick in purchases, primarily driven by lower prices during off-peak periods in the wholesale and retailer markets. The key takeaway is that as consumers become more willing to shift their consumption to periods of lower demand, the retailer has to buy from the wholesale market to fulfill demands during off-demand periods. This shift leads to an overall decrease in wholesale electricity prices and also enhances overall social welfare as detailed in Figure 7. From the perspective of market regulators, an important implication is that, depending on the market structure, tariffs-based incentives can have a significant impact on both spot and retail market outcomes.
@@ 4.5.1 Market outcomes: expected profit
Figure 6: Comparison of expected profit between two models before and after DR implementation.
The expected profit results of the retailer are depicted in Figure 6 in both models. The expected profit results are directly related to the price tariffs and quantities previously discussed. In the MILP model, the retailer’s myopic approach to market information results in financial losses in specific periods. This occurs because the model is specifically designed to maximize the expected total profit over an entire 24-hour time horizon [^3] The underlying assumption here is that consumers commit to a tariff agreement with the retailer, which is expected to remain valid for the specified time horizon. This approach is both reasonable and practical in the context of current electricity markets, as discussed in (Wu et al., 2015).. However, the consumers’ flexibility with DR plays a pivotal role, in enhancing the retailer’s expected profits and mitigating monetary losses. Notably, in the MPEC models, where the retailer adopts a strategic approach, its expected profits are significantly higher.
@@ 4.5.2 Market outcomes: Social welfare analysis
Figure 7: Social welfare for MPEC and equilibrium models in the No flexibility and Flexibility cases.
Figure 7 depicts social welfare with two models before and after the implementation of the DR program. Interestingly, the MPEC model explicitly highlights the social cost incurred when consumption is shifted to low-demand periods due to consumers’ flexibility. This suggests that price-maker retailers have the ability to manipulate this consumer flexibility, exploiting it to maximize their expected profits. In doing so, they transfer risks centered around the retailer to the consumers. On the other hand, it also demonstrates an increase in social welfare during peak-demand periods because consumers have the choice to adjust their consumption during these periods using dynamic price tariffs. Social welfare is better during low-demand periods due to low-price tariffs, and worse during peak-demand periods due to high-price tariffs before DR in both models. Despite the retailer’s exploitation, consumers’ flexibility is crucial in enhancing social welfare, especially during extended periods from 9:00 to 22:00. These hours are typically stressful for system operators, underlining the significance of consumer flexibility for system stability. In contrast, the equilibrium model exhibits positive social welfare in all periods. There are higher welfare gains during peak-demand periods, and the social cost during low-demand periods is significantly lower than that of the MPEC model. More importantly, with Figure 7(b), we present an analysis of social welfare before and after implementing the demand response program, using a Linear utility function. The findings from both models reveal that social welfare significantly improves with a linear utility function compared to a quadratic one. Interestingly, during off-peak periods, social welfare appears to be better without flexibility. However, despite the differences in the magnitude of welfare, the trend remains consistent whether using linear or quadratic utility functions. This consistency enables the generalization of market outcomes across the models, regardless of the specific functional form of the consumers’ utility functions.
@@
Figure 8 shows individual consumers’ welfare in both models, before and after implementing the DR program. Consumers generally experience higher surpluses, especially during low-demand periods by using MILP. Consumers’ flexibility leads to a surplus for consumers across all periods compared to their utility without the DR program. Even with the MPEC model, consumers offset their losses during low-demand in the time periods from 8:00 to 22:00. Note that the differences in individual consumers’ surplus are attributable to variations in utility and demand flexibility parameters, as detailed in Table 2.
Figure 9 presents results concerning consumers’ flexibility in their consumption for MPEC and MILP: the Benchmark and Flexibility cases. The overall trend for flexible demand is similar in both models, although there are differences in magnitude. Notably, consumers demonstrate increased flexibility from 9:00-11:00 and 17:00-22:00, shifting their consumption to off-peak periods. Thus, indicating higher energy consumption during these off-peak times. The results adhere to the constraint $\sum_{t\in{T}}\Delta_{jt\omega}^{C}=0$ and respect the flexibility boundary values for each consumer as presented in Table 2. For example, consumer 1 exhibits higher flexibility in consumption, up to 5kWh, compared to other consumers. With high flexibility, the overall electricity price decreases, and the quantity purchased increases, leading to robust impacts on both the retailer’s expected profit and consumer surplus. Therefore, the DR models effectively characterize the market outcomes in both market configurations, highlighting the importance of utility functions in understanding consumer behavior and market dynamics.
@@
Figure 9: Variation of demand in both models (left MPEC and right MILP) with respect to the Benchmark and Flexibility cases.
@@ 4.5.3 Market performance and sensitivity analysis
Table 4: Market performance of retailer and consumers in the simulations with the Benchmark case. All values, except for the price (€/kWh), represent average outcomes for the considered scenarios ($\Omega=30$) and are expressed in €.
@@
Table 4 presents the average performance of the two models for both the retailer and consumers under the Benchmark case. On average, the retailer incurs losses in the MILP model. The retailer has two sources of costs: i), cost incurred from purchasing in the spot market, which is based on the assumption of perfect information about future consumption; and ii) imbalance penalties that arise from imperfect information. It is essential to understand that the revenue generated by the retailer translates into costs for consumers, which significantly affects the overall dynamics of the market.
@@
In Table 5, we present the market outcomes regarding the uncertainty of spot market prices $P_{t\omega}^{S}$ under the Benchmark case and the MPEC model. On the contrary, consumers enjoy better utility due to the low price tariffs. We consider the coefficient of variations to be 0.015, 0.003, and 0.035. The results indicate that uncertainties in spot market prices do not significantly influence market outcomes for either retailers or consumers. Instead, as demonstrated by our simulation results, the most crucial parameter affecting market outcomes is demand flexibility. This flexibility in consumer demand is key not only in mitigating the potential impact on market outcomes but also in influencing spot market price formation. This highlights the pivotal role of consumer flexibility in managing market dynamics and aiding the integration of renewable energy sources.
@@ 5 Conclusions
This paper introduces two game-theoretical models to analyze retailer-consumer interactions in electricity markets under demand response programs. The paper model consumer flexibility under dynamic pricing and employs a utility function to characterize consumer behavior, providing insights into market dynamics and reshaping their interpretation.
@@
Two market structures are considered: a leader-follower model using a bi-level problem solved as MPEC and a perfect competition model solved as a MILP problem. The MPEC model facilitates a comparison of market power dynamics, while the MILP model helps in understanding competitive market behavior. Both models reveal their respective impacts on market outcomes and social welfare. This work advances existing research in demand response by introducing complex dynamics under dynamic price tariffs. It achieves this by simultaneously analyzing consumers’ utility functions, incorporating uncertainty, and considering various levels of market structures both in the spot and retail markets. These multifaceted interactions enable a more nuanced and comprehensive understanding of market dynamics, potentially reshaping how these dynamics are modeled and interpreted.
@@
The findings demonstrate the effectiveness of dynamic pricing in demand response programs, optimizing strategic retailers’ expected profit and maximizing consumers’ welfare. The models’ adaptability to various consumer groups and their flexibility to retailers’ pricing signals are crucial for demand-side management under demand response programs. Additionally, the research presents compelling economic incentives for flexible consumers under price uncertainties, benefiting both retailers and consumers and suggesting a symbiotic relationship between DR programs and electricity markets.
@@
The proposed methodology delivers crucial managerial and economic insights: i) it manages both demand and supply uncertainties through stochastic programming. ii) the analysis explores market configurations for each participant, focusing on enhancing flexibility in consumption and load management through dynamic tariffs. This approach aims to optimize the market structure for the overall improvement of social welfare. We have demonstrated that consumers’ flexibility is crucial in devising market strategies that are attuned to individual needs and adapt to varying market conditions. iii) the results reveal that utilizing consumers’ flexibility under tariff-based demand response renders significant cost savings and economic efficiency in electricity markets. The takeaway is that these benefits are crucial in promoting the sustainability and resilience of future electricity markets, underlying that demand response programs are key tools in energy transition under various market structures.
The proposed game-theoretical framework presents various opportunities for advanced research and application. First, it is possible to extend this framework to include the operational dynamics of generating companies and market operators. This enhancement would elucidate the reciprocal influences and impacts between these entities and demand response programs. Second, the models could be extended accounting for generation from renewable sources and grid constraints explicitly, coupled with the simulation of consumer behavior facilitated by smart appliances. This integration would enable the investigation of consumer-centric energy trading scenarios, evolving the role of consumers from passive end-users to proactive market contributors. Third, in light of the scant research on the environmental ramifications of demand response programs, integrating environmental metrics, particularly pollution metrics, into the demand response models is proposed. This could provide robust policy guidance for harmonizing technological advancements, market mechanisms, and innovative approaches within the electricity markets.

### v6

@@ 4.5 Market outcomes: Electricity prices and retailer’s quantities comparison
Now, let us compare price tariffs, quantities, consumers’ utility parameters, and demand flexibility to understand the dynamics of the market. Figure 5 depicts the price tariffs sent by the retailer to consumers and quantities purchased by the retailer from the spot market in both models. Price tariffs are plotted on the left axis, and spot market quantities are plotted on the right axis. Figure 5(a) depicts the benchmark case, which serves as a baseline for comparing the impacts of the market parameters on market outcomes. In the quadratic utility function, the linear term (as shown in Figure 5 (b)) captures the initial positive impact of electricity consumption on consumers’ utility. Moreover, the quadratic case models decrease in utility as electricity consumption increases. As a result, when the parameter $B_{jt\omega}$ increases, the retailer purchases a larger quantity from the wholesale market. This increase is illustrated in Figure 5(c). The retailer then sells this quantity to consumers at lower price tariffs in the retail market. A lower price, in turn, increases consumer utility. It should be noted that from the retailer’s perspective, an increase in $B_{jt\omega}$ has an effect analogous to an increase in demand from consumers.
Overall, the electricity price tariffs sent by the retailer and the quantities bought by the retailer from the spot market exhibit similar trends. In a competitive market, consumers’ utility increases due to lower prices. In the MPEC market, the expected profit of the retailer increases as it strategically affects price formation. While prices remain relatively stable in the MPEC model, the quantities purchased from the spot market are significantly affected by the flexibility of consumers. Most notably, greater flexibility has a greater impact on the quantity purchased in the spot market. A key takeaway is that as consumers become more willing to shift their consumption to periods of lower demand, the retailer has to buy from the wholesale market to fulfill demands during off-demand periods. This shift leads to an overall decrease in wholesale electricity prices and enhances overall social welfare as detailed in Figure 7. From these simple simulation results, the managerial insight is that market regulators should consider the market structure when designing an efficient demand response program to exploit the potential of consumer flexibility for both spot and retail market outcomes.
@@ 4.5.1 Market outcomes: expected profit
Figure 6: Comparison of the expected profit between the two models before and after DR implementation.
The expected profits of the retailer are depicted in Figure 6 for both models. They are directly related to price tariffs and quantities. If the retailer is myopic regarding the market, it loses money both overall and in some periods, as depicted by the MILP result. Note that in such models, the retailer maximizes its expected profit over an entire 24-hour time horizon, focusing more on the overall or cumulative profit than on profits from individual periods. [^3] The underlying assumption here is that consumers commit to a tariff agreement with the retailer, which is expected to remain valid for the specified time horizon. This approach is both reasonable and practical in the context of current electricity markets, as discussed in (Wu et al., 2015).. However, consumer flexibility with DR plays a pivotal role, in enhancing the retailer’s expected profits and mitigating financial losses. Notably, in the MPEC models, where the retailer is a strategic player, its expected profits are significantly greater due to its market power in price formation.
@@ 4.5.2 Market outcomes: Social welfare analysis
Figure 7: Social welfare for the MPEC and equilibrium models in the no flexibility and flexibility cases.
Figure 7 depicts social welfare with two models before and after the implementation of the DR program. The MPEC model shows that the social cost of consumption shifts to low-demand periods using consumers’ flexibility. This shows that retailers with market power can distort consumer flexibility, exploiting it to maximize their expected profits. In doing so, retailers with market power can transfer their risks to consumers. On the other hand, social welfare during peak-demand periods increases as consumers can adjust their consumption during these periods using dynamic price tariffs. Social welfare is better off during low-demand periods due to low-price tariffs, and worse off during peak-demand periods due to high-price tariffs before DR in both models. Consumer flexibility plays an important role in enhancing social welfare from 9:00 to 22:00. These hours are typically stressful for system operators, underlining the significance of consumer flexibility for system stability. Interestingly, the equilibrium model exhibits positive social welfare in all periods. Welfare is greater during peak-demand periods, and the social cost during low-demand periods is significantly lower than that of the MPEC model. Figure 7(b) presents social welfare before and after implementing the demand response program, using a linear utility function. The results from both models demonstrate that social welfare significantly improves with a linear utility function compared to a quadratic one. During off-peak periods, social welfare is better without consumer flexibility. However, despite differences in the magnitude of welfare, the trend remains consistent whether linear or quadratic utility functions are used. This consistency enables the generalization of market outcomes across the models, regardless of the specific functional form of the consumers’ utility functions.
@@
Figure 8 shows individual consumers’ welfare in both models, before and after implementing the DR program. Consumers generally experience surpluses, especially during low-demand periods when using MILP. Consumer flexibility leads to a surplus for consumers across all periods compared to their utility without the DR program. Even with the MPEC model, consumers offset their losses during low demand in the periods from 8:00 to 22:00. Note that the differences in individual consumer surpluses are attributable to variations in the utility and demand flexibility parameters, as detailed in Table 2.
Figure 9 presents consumer flexibility in the MPEC and MILP models for the benchmark and flexibility cases. The overall trend for flexibility is similar in both models. Notably, consumers demonstrate greater flexibility from 9:00-11:00 and 17:00-22:00 by shifting their consumption to off-peak periods. The results adhere to the constraint $\sum_{t\in{T}}\Delta_{jt\omega}^{C}=0$ and respect the flexibility boundary values for each consumer as presented in Table 2. For example, consumer 1 exhibits greater flexibility in consumption, up to 5 kWh, compared to other consumers. With high flexibility, the overall electricity price decreases, and the quantity purchased increases, leading to robust impacts on both the retailer’s expected profit and consumer surplus. Therefore, the DR models effectively characterize market outcomes under both market configurations. They highlight the importance of utility functions in understanding consumer behavior and market dynamics.
@@
Figure 9: Variation in demand in both models (left MPEC and right MILP) concerning the benchmark and flexibility cases.
@@ 4.5.3 Market performance and sensitivity analysis
Table 4: Market performance of the retailer and consumers in the simulations with the benchmark case. All values, except for the price (€/kWh), represent average outcomes for the considered scenarios ($\Omega=30$) and are expressed in €.
@@
Table 4 presents the average performance of the two models for both the retailer and consumers under the benchmark case. On average, the retailer incurs losses in the MILP model. The retailer’s costs come from two sources: i) the cost of purchasing in the spot market, which assumes perfect information about future consumption; and ii) imbalance penalties, which arise due to imperfect information. It is essential to understand that the revenue generated by the retailer translates into costs for consumers, which significantly affects the overall dynamics of the market.
@@
In Table 5, we present the market outcomes regarding the uncertainty of spot market prices $P_{t\omega}^{S}$ under the benchmark case and in the MPEC model. In contrast, consumers enjoy better utility due to low price tariffs. The results indicate that uncertainties in spot market prices do not significantly influence market outcomes for either retailers or consumers. As demonstrated by our simulation results, the most crucial parameter affecting market outcomes is consumer flexibility. Thus, consumer flexibility is key not only in mitigating the potential impact of the retailer’s market power on the market and in influencing spot market price formation. This highlights the pivotal role of consumer flexibility in managing market dynamics and aiding the integration of renewable energy sources.
@@ 5 Conclusions
This paper introduces two game-theoretical models to analyze retailer-consumer interactions in electricity markets under dynamic tariff-based demand response programs. It models consumer flexibility under dynamic pricing and employs a utility function to characterize consumer behavior, providing insights into market dynamics and reshaping their interpretation.
@@
Two market structures are considered: a leader-follower model using a bilevel problem solved as MPEC and a perfect competition model solved as a MILP problem. The MPEC model facilitates a comparison of market power dynamics, while the MILP model helps in understanding competitive market behavior. Both models reveal that the impact of consumer flexibility on market outcomes and social welfare depends on the specific market configuration. The paper advances existing research on demand response programs under dynamic price tariffs by incorporating uncertainty and levels of competition. It also explicitly accounts for consumer preferences through a utility function. These dynamic interactions allow for a more complete understanding of market dynamics and the potential of consumer flexibility. They further provide managerial insights for regulators on how to leverage this potential.
@@
The results demonstrate the effectiveness of dynamic pricing in demand response programs for optimizing strategic retailers’ expected profit and maximizing consumers’ welfare. It also presents economic incentives in the form of price reductions for flexible consumers under price uncertainties, which benefits both retailers and consumers.
@@
From the results, we can draw the following managerial and economic insights: i) the proposed models manage consumers’ demand and price uncertainties with stochastic programming, ii) the results show that market configurations play an important role in the effectiveness of consumer flexibility in demand response based on dynamic tariffs, iii) it is demonstrated that consumer flexibility could be crucial in devising market strategies that are attuned to individual needs and adapt to varying market conditions, and iv) the results reveal that utilizing consumer flexibility under tariff-based demand response renders significant cost savings and economic efficiency in day-ahead and retailer electricity markets. The takeaway is that enhancing consumer flexibility through demand response could play a crucial role in the sustainability and resilience of future electricity markets.
The proposed models can be extended in different directions. i) Including the operational dynamics of generating companies’ electricity generators and market operators might provide further insights into how to exploit the potential for consumer flexibility. ii) The models could be extended by explicitly accounting for global renewable requirements in operational decision-making and low-level smart grid constraints to facilitate the inclusion of prosumers in such market dynamics. iii) In light of the scant research on the environmental ramifications of demand response programs, integrating environmental metrics into demand response models is possible. This could provide robust policy guidance for harmonizing technological advancements, market mechanisms, and innovative approaches within electricity markets.